OpenStudy (anonymous):

Equation of the line parallel to the equation y = -2x - 7 that passes through the point (3, 1).

OpenStudy (anonymous):

y = -2x + 7 y = -2x + 6 y = -2x - 2 y = 3x - 8

OpenStudy (m.d.saxon):

y=-2x + 7

OpenStudy (mathstudent55):

Since the line is parallel, the slope has to be -2. That means you can eliminate the 4th choice. To decide between choices 1, 2, and 3, replace x and y in each choice with the x and y of the point you are given (x = 3, and y = 1), and see which choice works.

OpenStudy (m.d.saxon):

its the first one trut me I know